# Optimas: Optimizing Compound AI Systems with Globally Aligned Local Rewards

### v3

@@ 1 Introduction
Modern AI systems increasingly employ compound systems that integrate multiple complex components, such as Large Language Models (LLMs), tool/function calls, and traditional machine learning models like retrievers (Yuksekgonul et al., 2025; Khattab et al., 2023; Du et al., 2023). These components collaborate to process heterogeneous data sources and solve complex tasks through specialized subtask allocation (Zaharia et al., 2024; Zhou et al., 2025; Leike et al., 2018; Kandogan et al., 2025; Chen et al., 2025b). While compound AI systems have yielded performance advantages over monolithic models (Du et al., 2023; Huang et al., 2023; Leblond & et al., 2023; Lewis et al., 2020; Liu et al., 2023), they can be highly sensitive to the failure of individual components, which leads to cascading failures in the final results (Cemri et al., 2025; tse Huang et al., 2025; Peng et al., 2025; Chen et al., 2024a). For example, if an LLM misinterprets an input query, it can retrieve irrelevant or misleading information. This leads to subsequent tool calls operating on incorrect inputs, producing unreliable outputs throughout the system. Therefore, optimizing these systems as a whole is crucial for maintaining reliability and global system performance (i.e., global rewards).
@@
Present work. Here we propose Optimas (Figure 1), a unified framework for the effective and data-efficient optimization of compound AI systems. The core idea is to learn a globally aligned Local Reward Function (LRF) per component, such that independently maximizing a component’s local reward still reliably improves the global rewards. We show that under mild conditions, our approach converges reliably, providing strong theoretical guarantees. Furthermore, since the learned LRFs can be used to optimized components locally, Optimas has higher data efficiency by avoiding extensive runs of the entire compound AI system to achieve high global reward.
Specifically, each LRF estimates the contribution of a component’s output to the global reward. All LRFs are implemented using a shared LLM backbone, with separate projection heads for each component to produce component-specific rewards. We propose a lightweight adaptation mechanism using mini-batch preference data to ensure the LRFs remain aligned with the evolving system configuration (Figure 3). Leveraging this decentralized structure, Optimas applies specific optimization method to each component based on its configuration type. For example, reinforcement learning for model parameters (Rafailov et al., 2023; Schulman et al., 2017) or metric-guided search for prompts and hyperparameters (Yang et al., 2024; Opsahl-Ong et al., 2024; Liashchynskyi & Liashchynskyi, 2019). Overall, Optimas iteratively updates heterogeneous configurations towards a higher global reward by using each adaptive LRF as an objective. By optimizing a component to maximize its local reward, Optimas reduces the entire system runs to maintain higher data efficiency.
We conduct extensive experiments to evaluate Optimas across five real-world compound systems (Figure 2), including challenging settings such as behavior-driven product recommendation and medical analysis. Optimas consistently outperforms strong baselines, achieving an average relative improvement of 11.92% with higher data efficiency. While baseline methods occasionally improve performance. For example, while DSPy improves the performance on the multi-hop QA system, it may degrade performance on other tasks, such as product recommendation. In contrast, Optimas is the only method that improves performance across all five tasks, consistent with our theoretical guarantee (Section 4.4) that aligning local and global rewards enables effective optimization.
@@
Supports Optimizes Data efficiency Convergence guarantee
compound system heterogeneous config.
@@ 2 Related Work
Optimizing LLM single-step generation. Prior work extensively optimizes prompts for Large Language Models (LLMs) in single-step generation to improve performance (Yang et al., 2024; Madaan et al., 2023; Guo et al., 2024; Shinn et al., 2023; Yu et al., 2024; Yin et al., 2025; Chen et al., 2025a), but these methods are limited in their ability to handle complex, multi-step tasks. For example, addressing complex queries often requires combining multiple components—such as LLMs, tools, and machine learning predictors—to obtain accurate predictions.
@@ 4.1 Learning Local Reward Functions
where $\mathbf{v}_{-k}$ denotes the configurations of all downstream components (those that directly or indirectly receive information originating from $C_{k}$). The expected global reward for each candidate output is estimated via Monte Carlo sampling. This involves executing the downstream components with the candidate output fixed, capturing their stochasticity, and averaging the resulting global rewards from the final system outputs.
@@
The collection of $\mathcal{D}_{k}(\mathbf{v})$ follows the following steps: 1) execute the compound system up to $C_{k}$ and record the partial trajectory $\bigl\langle x,\,(x_{1},y_{1}),\dots,(x_{k-1},y_{\,k-1})\bigr\rangle$; 2) sample two candidate outputs for $C_{k}$ (e.g., via higher-temperature decoding or alternate hyperparameters); and 3) estimate their expected task metrics according to the expectation terms on the right-hand side of Eq. equation 3. The output with the higher expected value is labeled as $y_{k}^{+}$, and the other as $y_{k}^{-}$ in $\mathcal{D}_{k}(\mathbf{v})$.
@@ 4.2 Adaptive Local Reward Functions
Problem: Misaligned LRFs in the evolving system. As the system configuration changes during optimization, LRFs trained under a previous configuration $\mathbf{v}^{t}$ may become inaccurate under the updated configuration $\mathbf{v}^{t+1}$. Specifically, (1) after updating $C_{k}$, the same outputs from its upstream component $C_{i}$ ($i<k$) may lead to different global reward, making $r_{j}$ misaligned. (2) Its downstream components $C_{j}$ ($j>k$) now receive inputs generated by the updated $C_{k}$, which may fall outside the distribution seen by their LRFs. These shifts accumulate over time, degrading the local–global alignment property (Eq. 3) that LRFs are designed to satisfy.
@@
Stage 2: Online reward function adaptation. We sample a small batch of input data and construct a mini-batch of preference data $\mathcal{B}_{k}$ for $C_{k}$ using the steps described in Section 4.1. We then optimize the LRF on $C_{k}$ on the objective $\mathcal{L}_{k}(\mathcal{B}_{k})$ following the definition in Eq. 4. To enable stable optimization and improve data efficiency, we maintain a buffer of previous generated preference data into $\mathcal{B}_{k}$ . This adaptation helps maintain the local–global alignment property in Eq. 3.
@@ 4.3 Optimization with Globally aligned Local Reward Functions
Overall algorithm (Figure 3). Start with the initial system configuration, Optimas leverages initial reward modeling to learn a set of LRFs that are well-aligned with the global reward. At each iteration optimization, Optimas randomly selects a component to optimize, adapts its LRF to align with the evolving system using minimal amount of data, and conducts local optimization to obtain a new configuration. To prevent potential cascading errors, the new configuration is accepted only if it improves the system performance on a small validation set. Since the optimization with LRFs is conducted locally, the number of system runs to achieve a high global reward is reduced, as we later show in the experiments. The detailed algorithm is provided in Appendix A.
@@ 4.4 Theoretical Insights
Under regulartiy conditions, the maximizer of equation 4 satisfies the local-global alignment property (equation 3). In addition, maximizing $r_{k}\bigl(x_{k},\,C_{k}(x_{k};\mathbf{v}_{k})\bigr)$ over $\mathbf{v}_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $\mathbf{v}_{k}$ will yield the same solution.
@@
In fact, our theoretical analysis shows that by conducting local optimization, Optimas is essentially performing coordinate maximization. Therefore, existing convergence results for coordinate maximization directly apply. For instance, if the objective function additionally satisfies the Polyak–Łojasiewicz (PL) condition, the algorithm is guaranteed to converge to the global maximum.
@@ 5 Experiments
  • PubMedQA (Jin et al., 2019): A clinical classification dataset derived from PubMed abstracts (National Center for Biotechnology Information (2024), NCBI), evaluated by accuracy, defined as the proportion of predictions that exactly match the ground-truth labels.
@@
Baselines (See Appendix E for details). We compare Optimas against five baselines: Unoptimized, LLMSelector (Chen et al., 2025b), Hierarchical Behavior Cloning (HBC) (Le et al., 2018), TextGrad (Yuksekgonul et al., 2025), DSPy (Khattab et al., 2023; Opsahl-Ong et al., 2024).
@@
While the strong baseline DSPy shows notable improvements on some datasets (e.g., a 21.6% gain on HotpotQA), their performance may be inconsistent and can even degrade the system (e.g., a 14.3% drop on the Amazon dataset). For LLMSelector, it requires $2.8k$ times of forwarding through the entire system, which is $3x$ more expensive than Optimas.
@@ 5.2 Why and How Optimas Works: Alignment, Interpretability, and Efficiency
Amazon PubMedQA STaRK-Prime HotpotQA BigCodeBench Avg.
LLM Judge (20-shot) 51.25% 49.54% 54.37% 50.00% 42.45% 49.52%
Optimas 84.93% 65.28% 76.64% 72.40% 90.57% 77.96%
Takeaway 3: Optimas yields high local-global alignment quality (Table 4). To measure alignment quality of LRFs, we compute pairwise ranking accuracy: the probability that an output with higher global reward receives a higher score than an output with lower global reward. This reflects how well the learned LRFs aligns with global rewards. We compare against a LLM Judge, which prompts a gpt-4o model to score the outputs of components based on 20 in-context examples. This approach is similar to prior methods such as TextGrad, which rely on few-shot reasoning over textual patterns. In Table 4, LLM Judge performs closer to random guessing, due to the diversity and stochasticity of components’s outputs that make it difficult to reason reliably. In contrast, our LRFs achieve substantially higher performances. Moreover, LRFs internalize the local-global alignment within their weights without relying on limited in-context examples, enabling more precise alignment.
@@
Takeaway 6: Optimas is data and computational efficient. (Appendix G). We also conduct extensive studies to show the data efficiency with varying dataset size and computational efficiency with lightweight models. Due to space limitation, please refer to Appendix G.
@@ 6 Conclusion
Ethics statement.
@@
We provide full open access to our implementation code, including the original datasets, compound system implementations, and the Optimas framework at https://anonymous.4open.science/r/optimas-1308/. To support reproducibility, Appendix C (Dataset Details), Appendix D (System Details), Appendix E (Baseline Details), and finally, Appendix F (Optimas Details) further documents the datasets used and model architectures etc.. Together with our descriptions in the experiment section, we ensure that our data and code are fully accessible and our results are transparent and easily extensible.
@@ Appendix A Algorithm
13:Reward-model adaptation
14:  $D_{t}\leftarrow\textsc{CollectPreferenceData}(\mathcal{D},\Theta^{t},i_{t},k)$ $\triangleright$ create $(x^{+},x^{-})$ pairs for $C_{i_{t}}$ based on $R(\cdot)$
15:  $\theta^{t+1}_{i_{t}}\leftarrow\textsc{RewardModelTrain}(\theta^{t}_{i_{t}},D_{t})$
16:  $\Theta^{t+1}\leftarrow\Theta^{t}$ with $\theta_{i_{t}}$ replaced by $\theta^{t+1}_{i_{t}}$
17:Local optimization for the chosen component $C_{i_{t}}$
18:  if $C_{i_{t}}$ is an LLM with prompts then
19:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PromptOptimization}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
20:  else if $C_{i_{t}}$ has trainable weights then
21:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PPOTrain}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
22:  else if $C_{i_{t}}$ has a hyperparameter configuration then
23:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{HyperparameterSearch}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
24:  end if
25:  $\tilde{\mathbf{v}}\leftarrow\mathbf{v}^{t}$ with $\mathbf{v}^{t}_{i_{t}}$ replaced by $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}$
26:Validation
27:  if $\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\tilde{\mathbf{v}})\big)\;>\;\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\mathbf{v}^{t})\big)$ then
28:   $\mathbf{v}^{t+1}\leftarrow\tilde{\mathbf{v}}$;  $\mathbf{v}^{*}\leftarrow\mathbf{v}^{t+1}$
29:  else
30:   $\mathbf{v}^{t+1}\leftarrow\mathbf{v}^{t}$

### v4

@@ 1 Introduction
Modern AI systems increasingly employ compound systems that integrate multiple complex components, such as Large Language Models (LLMs), tool/function calls, and traditional machine learning models like retrievers (Yuksekgonul et al., 2025; Khattab et al., 2023; Du et al., 2023). These components collaborate to process heterogeneous data sources and solve complex tasks through specialized subtask allocation (Zaharia et al., 2024; Zhou et al., 2025; Leike et al., 2018; Kandogan et al., 2025; Chen et al., 2025b). While compound AI systems have yielded performance advantages over monolithic models (Du et al., 2023; Huang et al., 2023; Leblond and et al., 2023; Lewis et al., 2020; Liu et al., 2023), they can be highly sensitive to the failure of individual components, which leads to cascading failures in the final results (Cemri et al., 2025; Huang et al., 2025; Peng et al., 2025; Chen et al., 2024a). For example, if an LLM misinterprets an input query, it can retrieve irrelevant or misleading information. This leads to subsequent tool calls operating on incorrect inputs, producing unreliable outputs throughout the system. Therefore, optimizing these systems as a whole is crucial for maintaining reliability and global system performance (i.e., global rewards).
@@
Present work. Here we propose Optimas (Figure 1), a unified framework for the effective and data-efficient optimization of compound AI systems. The core idea is to learn a globally aligned Local Reward Function (LRF) per component, such that independently maximizing a component’s local reward still reliably improves the global rewards. We show that under mild conditions, our approach converges reliably, providing strong theoretical guarantees. Furthermore, since the learned LRFs can be used to optimize components locally, Optimas has higher data efficiency by avoiding extensive runs of the entire compound AI system to achieve high global reward.
Specifically, each LRF estimates the contribution of a component’s output to the global reward. All LRFs are implemented using a shared LLM backbone, with separate projection heads for each component to produce component-specific rewards. We propose a lightweight adaptation mechanism using mini-batch preference data to ensure the LRFs remain aligned with the evolving system configuration (Figure 3). Leveraging this decentralized structure, Optimas applies specific optimization method to each component based on its configuration type. For example, reinforcement learning for model parameters (Rafailov et al., 2023; Schulman et al., 2017) or metric-guided search for prompts and hyperparameters (Yang et al., 2024; Opsahl-Ong et al., 2024; Liashchynskyi and Liashchynskyi, 2019). Overall, Optimas iteratively updates heterogeneous configurations towards a higher global reward by using each adaptive LRF as an objective. By optimizing a component to maximize its local reward, Optimas reduces the entire system runs to maintain higher data efficiency.
We conduct extensive experiments to evaluate Optimas across five real-world compound systems (Figure 2), including challenging settings such as behavior-driven product recommendation and medical analysis. Optimas consistently outperforms strong baselines, achieving an average relative improvement of 11.92% with higher data efficiency, while baseline methods occasionally improve performance. For example, while DSPy improves the performance on the multi-hop QA system, it may degrade performance on other tasks, such as product recommendation. In contrast, Optimas is the only method that improves performance across all five tasks, consistent with our theoretical guarantee (Section 4.4) that aligning local and global rewards enables effective optimization.
@@
Supports Optimizes Data Convergence
compound system heterogeneous config. efficiency guarantee
@@ 2 Related Work
Optimizing LLM single-step generation. Prior work extensively optimizes prompts for Large Language Models (LLMs) in single-step generation to improve performance (Yang et al., 2024; Madaan et al., 2023; Guo et al., 2024; Shinn et al., 2023; Yu et al., 2024; Yin et al., 2025; Chen et al., 2025a; Williams, 1992), but these methods are limited in their ability to handle complex, multi-step tasks. For example, addressing complex queries often requires combining multiple components—such as LLMs, tools, and machine learning predictors—to obtain accurate predictions.
@@ 4.1 Learning Local Reward Functions
where $\mathbf{v}_{-k}$ denotes the configurations of all downstream components (those that directly or indirectly receive information originating from $C_{k}$). The expected global reward for each candidate output is estimated via Monte Carlo sampling. This involves executing the downstream components with the candidate output and the outputs from the non-downstream components fixed, capturing their stochasticity, and averaging the resulting global rewards from the final system outputs.
@@
The collection of $\mathcal{D}_{k}(\mathbf{v})$ follows the following steps: 1) execute the compound system up to $C_{k}$ and record the partial trajectory $\bigl\langle x,\,(x_{1},y_{1}),\dots,(x_{k-1},y_{\,k-1})\bigr\rangle$; 2) sample two candidate outputs for $C_{k}$ (e.g., via higher-temperature decoding or alternate hyperparameters); and 3) estimate their expected task metrics according to the expectation terms on the right-hand side of Eq. 3. The output with the higher expected value is labeled as $y_{k}^{+}$, and the other as $y_{k}^{-}$ in $\mathcal{D}_{k}(\mathbf{v})$.
@@ 4.2 Adaptive Local Reward Functions
Problem: Misaligned LRFs in the evolving system. As the system configuration changes during optimization, LRFs trained under a previous configuration $\mathbf{v}^{t}$ may become inaccurate under the updated configuration $\mathbf{v}^{t+1}$. Specifically, (1) after updating $C_{k}$, the same outputs from its upstream component $C_{i}$ ($i<k$) may lead to different global reward, making $r_{i}$ misaligned. (2) Its downstream components $C_{j}$ ($j>k$) now receive inputs generated by the updated $C_{k}$, which may fall outside the distribution seen by their LRFs. These shifts accumulate over time, degrading the local–global alignment property (Eq. 3) that LRFs are designed to satisfy.
@@
Stage 2: Online reward function adaptation. When any configuration changes, we sample a small batch of input data and construct a mini-batch of preference data $\mathcal{B}_{k}$ for each component $C_{k}$ using the steps described in Section 4.1. We then optimize the LRF on $C_{k}$ on the objective $\mathcal{L}_{k}(\mathcal{B}_{k})$ following the definition in Eq. 4. To enable stable optimization and improve data efficiency, we maintain a buffer of previous generated preference data into $\mathcal{B}_{k}$ . This adaptation helps maintain the local–global alignment property in Eq. 3.
@@ 4.3 Optimization with Globally aligned Local Reward Functions
Overall algorithm (Figure 3). Starting with the initial system configuration, Optimas leverages initial reward modeling to learn a set of LRFs that are well-aligned with the global reward. At each iteration optimization, Optimas randomly selects a component to optimize, conducts local optimization, and if the local configuration change leads to improved global reward, it updates the system and adapts the LRF using minimal amount of data. To prevent potential cascading errors, the new configuration is accepted only if it improves the global reward on a small validation set. Since the optimization with LRFs is conducted locally, the number of system runs to achieve a high global reward is reduced, as we later show in the experiments. The detailed algorithm is provided in Appendix A.
@@ 4.4 Theoretical Insights
Under regularity conditions, the minimizer of equation 4 satisfies the local-global alignment property (equation 3). In addition, maximizing $r_{k}\bigl(x_{k},\,C_{k}(x_{k};\mathbf{v}_{k})\bigr)$ over $\mathbf{v}_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $\mathbf{v}_{k}$ will yield the same solution.
@@
In fact, our theoretical analysis shows that by conducting local optimization, Optimas is essentially performing coordinate maximization. Therefore, existing convergence results for coordinate maximization directly apply. Also, note that the block-coordinate (round-robin) updates adopted in Optimas do not guarantee global optimality in non-convex problems, but this does not constitute a flaw unique to our method; rather, it reflects a standard limitation of non-convex optimization broadly, and our global convergence guarantees only hold under additional structural assumptions, such as Polyak–Łojasiewicz or Kurdyka–Łojasiewicz conditions.
@@ 5 Experiments
  • PubMedQA (Jin et al., 2019): A clinical classification dataset derived from PubMed abstracts (National Center for Biotechnology Information (NCBI), 2024), evaluated by accuracy, defined as the proportion of predictions that exactly match the ground-truth labels.
@@
Baselines (See Appendix E for details). We compare Optimas against five baselines: Unoptimized, REINFORCE (Williams, 1992), LLMSelector (Chen et al., 2025b), Hierarchical Behavior Cloning (HBC) (Le et al., 2018), TextGrad (Yuksekgonul et al., 2025), DSPy (Khattab et al., 2023; Opsahl-Ong et al., 2024). Moreover, we provide a single LLM reference which prompts an LLM to complete the task directly. This reference is used for justifying our system design in the experimental setup.
@@
Amazon PubMedQA STaRK-Prime HotpotQA BigCodeBench Avg.
LLM Judge 51.25% 49.54% 54.37% 50.00% 42.45% 49.52%
Optimas 84.93% 65.28% 76.64% 72.40% 90.57% 77.96%
@@
For REINFORCE, while it yields performance gain from the unoptimized baseline, Optimas outperforms it by around 3%. Moreover, REINFORCE needs collecting the reward signal from downstream Monte Carlo sampling, which requires more than three times the data than Optimas. While the strong baseline DSPy shows notable improvements on some datasets (e.g., a 21.6% gain on HotpotQA), their performance may be inconsistent and can even degrade the system (e.g., a 14.3% drop on the Amazon dataset). For LLMSelector, it requires $2.8k$ times of forwarding through the entire system, which is $3x$ more expensive than Optimas.
@@ 5.2 Why and How Optimas Works: Alignment, Interpretability, and Efficiency
Takeaway 3: Optimas yields high local-global alignment quality (Table 4). To measure alignment quality of LRFs, we compute pairwise ranking accuracy: the probability that an output with higher global reward receives a higher score than an output with lower global reward. This reflects how well the learned LRFs aligns with global rewards. We compare against a LLM Judge, which prompts a gpt-4o model to score the outputs of components based on 20 in-context examples. This approach is similar to prior methods such as TextGrad, which rely on few-shot reasoning over textual patterns. In Table 4, LLM Judge performs closer to random guessing, due to the diversity and stochasticity of components’ outputs that make it difficult to reason reliably. In contrast, our LRFs achieve substantially higher performances. Moreover, LRFs internalize the local-global alignment within their weights without relying on limited in-context examples, enabling more precise alignment.
@@
Backbone Alignment
size quality (%)
1B 72.58
3B 70.04
8B 71.20
Figure 7: Impact of training data size and LRF backbone size on alignment quality, measured by pairwise ranking accuracy: the percentage of preference pairs where local and global rewards agree.
Takeaway 6: Optimas is data- and computationally efficient. (Figure 7 & Appendix G). We study the efficiency of Optimas along two axes on the HotpotQA system: (1) how much training data is needed to learn effective LRFs, and (2) whether large LRF backbones are necessary.
We train LRF models using varying percentages of available training data (12.5%–100%) and measure alignment quality via pairwise ranking accuracy. Figure 7 (left) shows the performance degrades gracefully: using just 12.5% of the data yields 65.46% accuracy (92.7% of full-data performance), and 25% recovers 95.1%. This indicates that Optimas is data efficient, where LRF can be learned from relatively few system runs. Moreover, Figure 7 (right) compares 1B, 3B, and 8B LRF backbones and reports alignment quality. The results show that Optimas is computationally efficient, where lightweight models are sufficient for learning local-global alignment.
In Appendix G, we further show that local and global reward landscapes are closely aligned across retriever top-$k$ settings, that modest numbers of prompt candidates and adaptation inputs suffice for strong performance, and provide a cost comparison on the Amazon system showing that Optimas achieves the best performance while using a comparable number of effective system runs.
@@ 6 Conclusion
Acknowledgements
We gratefully acknowledge the support of Amazon, NSF under Nos. CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); NIH under No. 1U24NS146314-01, Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Genentech, SAP, and SCBX.
Ethics Statement
@@ Acknowledgements
We provide full open access to our implementation code, including the original datasets, compound system implementations, and the Optimas framework. To support reproducibility, we provide Appendix C (Dataset Details), Appendix D (System Details), Appendix E (Baseline Details), and finally, Appendix F (Optimas Details).
@@ Appendix A Algorithm
13:Local optimization for the chosen component $C_{i_{t}}$
14:  if $C_{i_{t}}$ is an LLM with prompts then
15:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PromptOptimization}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
16:  else if $C_{i_{t}}$ has trainable weights then
17:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PPOTrain}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
18:  else if $C_{i_{t}}$ has a hyperparameter configuration then
19:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{HyperparameterSearch}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
20:  end if
21:  $\tilde{\mathbf{v}}\leftarrow\mathbf{v}^{t}$ with $\mathbf{v}^{t}_{i_{t}}$ replaced by $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}$
22:Validation
23:  if $\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\tilde{\mathbf{v}})\big)\;>\;\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\mathbf{v}^{t})\big)$ then
24:   $\mathbf{v}^{t+1}\leftarrow\tilde{\mathbf{v}}$;  $\mathbf{v}^{*}\leftarrow\mathbf{v}^{t+1}$
25:   Reward-model adaptation
26:   $D_{t}\leftarrow\textsc{CollectPreferenceData}(\mathcal{D},\Theta^{t},k)$ $\triangleright$ create $k$ $(x^{+},x^{-})$ pairs
27:   $\Theta^{t+1}\leftarrow\textsc{RewardModelTrain}(\Theta^{t},D_{t})$
28:  else
29:   $\mathbf{v}^{t+1}\leftarrow\mathbf{v}^{t}$
30:   $\mathbf{\Theta}^{t+1}\leftarrow\mathbf{\Theta}^{t}$